It means that there are at most five times in which the sought event occurs.
It means that there are at most five times in which the sought event occurs.
It means that there are at most five times in which the sought event occurs.
It means that there are at most five times in which the sought event occurs.
What are the odds of rolling 6 consecutive doubles with 2 dice?
The odds of rolling 6 out of 2 consecutive dice is 100/25.
Stringent means tightly controlled, detailed, harsh. It often carries a connotation of suspicion, particularly when used to describe looking into a person's background.
How many permutations are there of the latters of the word numbers?
There are 7 factorial, or 5040 permutations of the letters in the word NUMBERS.
In a standard deck of cards there is one and only one ace of diamonds.
The probability is that the remaining two outcomes are 1 H and 1 T. The probability of that is 2/4 or 50%
The probability is that the remaining two outcomes are 1 H and 1 T. The probability of that is 2/4 or 50%
The probability is that the remaining two outcomes are 1 H and 1 T. The probability of that is 2/4 or 50%
The probability is that the remaining two outcomes are 1 H and 1 T. The probability of that is 2/4 or 50%
Why the normal distribution curve have extremes is opened?
Because the domain of the normal distribution is infinite - in both directions.
How many ways arrange the word statistics?
The letters of the word statistics can be permuted in 10!/(3!*3!*2!) = 3,628,800/72 = 50,400 ways.
If there are 117 blue marbles229 red marbles and 104 yellow marbleswhat percent are blue?
26 percent because
117
229
+104
_____
450 ÷ 45 =10 x 10 =100
117 ÷ 45 =2.6 x 10 = 26
26
is = to 26%
100
Alright these are two different independent trials. Therefore u take the probabilityof the first times the probability of the second so 3/6 times 3/6 and u get 9/36=1/4=.25=25%
I rolled the dice 120 times what's the probability of getting a 6?
If you rolled a die 120 times, the probability of getting a 6 is one in six. It does not matter how many times you roll the die - the probability is still one in six - except that the long term mean will approach the theoretical value of 0.166... as the number of trials increases.
1/2
What is the probability of scoring a four on a die?
If you throw a fair die one time, the chance of a four is 1/6 or about .1667
What is the probability of drawing an Ace or a Jack from a deck of 52 cards?
Reason: There are 2 cards in each suit of 13 that are either an Ace or a Jack.
So there is a 2 out of 13 (or 8 out of 52) chance, which is 15.38% chance.
IN FRACTION AND USING PERMUTATIONS AND COMBINATIONStotal number of cases : 52C1
favorable cases : in cards there are 4 aces and 4 jacks, therefore 4+4 = 8
P (of drawing an ace or a jack) = 8/52
= 2/13
odds against this question is : 2/(13-2)
= 2/11 [ANS.]
probably means that something or which is not sure .
LIKE :- I PROBABLY GET THIS ANSWER RIGHT.